8x =4x(squared) -2
Answer by stanbon(75887)

8x =4x(squared) -2
8x = 4x^2 -2
4x^2 - 8x -2 = 0
2x^2 - 4x -1 = 0
Use the quadratic formula to get:
x = [4 +- sqrt(16 -4*2*-1)]/4
x = [4 +- sqrt(24)]/4
x = [4+-2sqrt(6)]/4
x = [1 + (1/2)sqrt(6)] or x = [1 - (1/2)sqrt(6)]
==============
Cheers,
Stan H.